Solve this equation using any method.
x+y-z=-20
2x-y+z=11

Answer by nyc_function(2741)

After playing with this strange question, I came up with the following:
Add the two equations and you get:
3x = -9 so x = -3
Now substitute this value to get:
y -z = -17
-y +z = 17
So, y = -17 + z for any value of z. ie. infinite number of solutions.
Try it out.